OpenStudy (anonymous):

The height h in feet of a ball after t seconds is given by h=128t-16t^2. How long did it take for the ball to hit the ground?

OpenStudy (anonymous):

h=128t-16t^2 height=0 ,when ball hits the ground 128t-16t^2=0 16t(8-t)=0 either t=0 or (8-t)=0 t=8 answer: t=0, t=8 hope this is the way.....sorry if I went wrong.